# Observer-robust energy condition verification for warp drive spacetimes

### v1

@@ 8 Discussion
The central result of this work is that single-frame energy condition evaluation can systematically underestimate the extent of energy condition violations in warp drive spacetimes, though the severity depends strongly on the metric geometry. For the Rodal metric, the Eulerian DEC analysis fails to detect violations at over 28% of all grid points, and the WEC figure exceeds 15%. For WarpShell (regularized implementation), missed WEC violations are below $0.1\%$ of the grid. For the Alcubierre and Natário metrics, we find no additional NEC or WEC violating grid points beyond the baseline Eulerian analysis at the tested parameters (missed =0=0 at the sampled resolution). Lentz shows zero missed NEC/WEC points (though nonzero SEC misses exist) and is severely under-resolved ($\ll\!1$ cell; Table 2); this result should be treated as a lower bound. Van Den Broeck exhibits small but nonzero missed fractions (NEC $0.1\%$, WEC $0.4\%$, DEC $0.3\%$ at $v_{s}=0.5$), and non-trivial SEC misses persist, most notably $1.2\%$ for Van Den Broeck (Table 3). Even for these metrics the observer-optimized violation severity can be much larger (Section 2.4, Interpretation caveat): for the Alcubierre WEC at $\zeta_{\max}=5$, the capped minimum is $\sim\!-3400$ compared to 0.038-0.038 for the Eulerian frame. These severity ratios are $\zeta_{\max}$-dependent and grow without bound; single-frame analysis can underestimate the magnitude of violations even when it correctly identifies their spatial extent.
@@
For the Alcubierre and Natário metrics, the baseline analysis detects all NEC violation points; Lentz also shows zero missed NEC points but is severely under-resolved ($\ll\!1$ cell; Table 2). Van Den Broeck shows small residual NEC/WEC misses ($0.1\%$$0.4\%$), and SEC misses remain (up to $1.2\%$ for Van Den Broeck). The worst-case WEC severity for boosted observers can be much larger (Section 2.4, Interpretation caveat); the specific ratios are $\zeta_{\max}$-dependent. The violation-set agreement may reflect the high symmetry of these metrics, where the ADM normal is naturally aligned with the principal stress-energy eigenvectors. We hypothesize that the Rodal metric breaks this alignment with its irrotational angular shift component, which would explain both the additional violation points and severity increase; the observer-alignment analysis (Figure 12) confirms this misalignment quantitatively.
Figure 12: Histogram of the alignment angle $\theta=\arccos|\hat{s}\cdot\hat{e}_{\text{spatial}}|$ between the optimizer boost direction $\hat{s}$ (unit spatial 3-vector of the worst-case observer’s velocity) and the spatial part $\hat{e}_{\text{spatial}}$ of the DEC-worst eigenvector of $T^{a}{}_{b}$ (the eigenvector whose principal pressure $|p_{i}|$ most exceeds $\rho$), for the Rodal metric ($50^{3}$ grid, $\zeta_{\max}=5$; see Table 1). Both vectors are normalized in the Euclidean (coordinate) norm; the absolute value folds the eigenvector sign ambiguity into $0^{\circ},90^{\circ}]$. Each panel shows $\sim\!108\,000$ Type I DEC-violating points at the indicated $v_{s}$. Medians range from $76^{\circ}$ ($v_{s}=0.1$) to $85^{\circ}$ ($v_{s}=0.99$); across all four velocities the combined median is $82^{\circ}$ (near-orthogonal), confirming that the optimizer boost direction is far from the eigenvector prediction and that single-frame analysis misses violations because the Eulerian frame does not probe these directions. Only 0.20.21.21.2% of points have alignment below $10^{\circ}$.
@@
Since the vast majority of grid points classify as Hawking–Ellis Type I (Table 8), the eigenvalue criteria $\rho\geq 0$, $\rho+p_{i}\geq 0$, and $\rho\geq|p_{i}|$ are the primary determinant of energy condition satisfaction or violation at most points. For Type I, these algebraic slacks provide exact certification: a Type I point satisfies an energy condition if and only if the corresponding eigenvalue inequality holds, regardless of observer, and the result is $\zeta_{\max}$-independent. The observer optimization serves two additional roles: it identifies the worst-case observer (boost direction and rapidity), and it provides a cross-check against the algebraic result. For non-Type-I stress-energy tensors, which do not admit a diagonal eigenvalue representation, the optimization becomes the sole method for evaluating observer-dependent conditions, and the algebraic shortcut does not apply. At such points, the reported WEC/SEC/DEC margins are $\zeta_{\max}$-capped lower bounds: a negative value certifies a genuine violation, but a positive value indicates only that no violation was found within the rapidity cap, not that the condition is satisfied for all observers. We validate this pathway with a null-dust (Type II) benchmark in the test suite (see Section 8).
@@
The observer optimizer with $N_{\text{starts}}=8$ random restarts is not guaranteed to find the global minimum; for highly non-convex objectives, some local minima may be missed. The soft rapidity cap $\zeta_{\max}\tanh(|\mathbf{w}|/\zeta_{\max})$ attenuates gradients near the rapidity bound, which could impede finding near-lightlike worst-case observers. Concretely, the $\tanh$ map has derivative $\mathrm{sech}^{2}(|\mathbf{w}|/\zeta_{\max})$, which decays exponentially as $|\mathbf{w}|\to\infty$; an optimizer step near the rapidity ceiling therefore receives a vanishingly small gradient signal, and worst-case observers that require $\zeta$ close to $\zeta_{\max}$ may not be found to full precision. A constrained optimizer (projected gradient or L-BFGS-B with hard bounds) would eliminate this failure mode and is planned for a future release. The boost-vector normalization uses the smooth approximation $|\mathbf{w}|_{\varepsilon}=\sqrt{\mathbf{w}\cdot\mathbf{w}+\varepsilon^{2}}$ (with $\varepsilon=10^{-12}$) to remain $C^{\infty}$ at $\mathbf{w}=0$, avoiding discontinuous gradients that could confuse quasi-Newton Hessian updates near the Eulerian point. The ablation study in Section 7.2 confirms that margins and missed-violation fractions are effectively converged by $N_{\text{starts}}=4$, with $N_{\text{starts}}=8$ providing a safety margin.
Since most warp metrics in our test suite are predominantly Hawking–Ellis Type I, the observer optimizer is not extensively demonstrated on a regime where it is strictly necessary for violation detection (algebraic checks suffice for Type I). To validate the non-Type-I pathway, we include a null-dust benchmark (Type II by construction, $T_{ab}=\Phi^{2}k_{a}k_{b}$) in the test suite, verifying that the classifier correctly identifies the null eigenvector structure and the optimizer recovers the expected NEC saturation ($T_{ab}k^{a}k^{b}=0$ to $10^{-12}$ residual) and non-negative WEC ($T_{ab}u^{a}u^{b}\geq 0$) for all tested observers; full numerical results are in the repository test suite. Full Richardson extrapolation (Section 7) was performed for the Alcubierre metric; the three-point data is non-monotone, so the assumed second-order extrapolation confirms resolution stability rather than a measured convergence rate. Rodal and WarpShell were assessed via resolution stability checks (missed-fraction and margin variation across $25^{3}$, $50^{3}$, and $100^{3}$ grids) but were not subjected to formal Richardson extrapolation, so the reported violation fractions for those metrics carry additional resolution uncertainty.
@@
  • Shell boundaries ($R_{1}=0.5$, $R_{2}=1.0$): centers of the Hermite-smoothstep shell indicator ($C^{2}$ smooth) that controls the spatial-metric conformal factor.
@@ 9 Conclusion
    Violation severity exceeds single-frame estimates. For the Alcubierre and Natário metrics, the baseline analysis detects all NEC violation points; Lentz shows zero missed NEC/WEC points (nonzero SEC misses exist) and is severely under-resolved ($\ll\!1$ cell; this result is a lower bound). Van Den Broeck exhibits small residual NEC/WEC misses ($0.1\%$$0.4\%$ at $v_{s}=0.5$), and SEC misses persist (up to $1.2\%$ for Van Den Broeck). The observer-optimized WEC severity at $\zeta_{\max}=5$ is much larger (e.g. Alcubierre WEC: $\sim\!90{,}000\times$; this ratio is $\zeta_{\max}$-dependent and grows as $\sim\!e^{2\zeta_{\max}}$). A rapidity-cap sensitivity study (Table 9) confirms that the robust WEC margin scales as $\sim\!e^{2\zeta_{\max}}$, while the NEC margin and violation-point set are cap-independent.
@@
    Tidal forces at bubble walls are significant. Geodesic deviation through the Alcubierre bubble shows sharp tidal eigenvalue spikes at the bubble wall, consistent with the expectation that steeper walls increase curvature gradients. The peak blueshift at the bubble wall reaches $7.1\times$ at $v_{s}=0.99$ compared to $v_{s}=0.1$.
@@
The author acknowledges support from the VinUni Center for Environmental Intelligence (CEI).

### v2

@@ 8 Discussion
The central result of this work is that single-frame energy condition evaluation can systematically underestimate the extent of energy condition violations in warp drive spacetimes, though the severity depends strongly on the metric geometry. For the Rodal metric, the Eulerian DEC analysis fails to detect violations at over 28% of all grid points, and the WEC figure exceeds 15%. For WarpShell (regularized implementation), missed WEC violations are below $0.1\%$ of the grid. For the Alcubierre and Natário metrics, we find no additional NEC or WEC violating grid points beyond the baseline Eulerian analysis at the tested parameters (missed =0=0 at the sampled resolution). Lentz shows zero missed NEC/WEC points (though nonzero SEC misses exist) and is severely under-resolved ($\ll\!1$ cell; Table 2); this result should be treated as a lower bound. Van Den Broeck exhibits small but nonzero missed fractions (NEC $0.1\%$, WEC $0.4\%$, DEC $0.3\%$ at $v_{s}=0.5$; Figures 34), and non-trivial SEC misses persist, most notably $1.2\%$ for Van Den Broeck (conditional miss rate $17.5\%$; Table 3). Even for these metrics the $\zeta_{\max}$-capped violation severity can be much larger (Section 2.4, Interpretation caveat): for the Alcubierre WEC at $\zeta_{\max}=5$, the capped minimum is $\sim\!-3450$ compared to 0.038-0.038 for the Eulerian frame. These severity ratios are $\zeta_{\max}$-dependent and grow without bound; single-frame analysis can underestimate the magnitude of violations even when it correctly identifies their spatial extent.
@@
For the Alcubierre and Natário metrics, the baseline analysis detects all NEC violation points; Lentz also shows zero missed NEC points but is severely under-resolved ($\ll\!1$ cell; Table 2). Van Den Broeck shows small residual NEC/WEC misses ($0.1\%$$0.4\%$; Figures 34), and SEC misses remain (up to $1.2\%$ for Van Den Broeck, conditional miss rate $17.5\%$). The worst-case WEC severity for boosted observers can be much larger (Section 2.4, Interpretation caveat); the specific ratios are $\zeta_{\max}$-dependent. The violation-set agreement may reflect the high symmetry of these metrics, where the ADM normal is naturally aligned with the principal stress-energy eigenvectors. We hypothesize that the Rodal metric breaks this alignment with its irrotational angular shift component, which would explain both the additional violation points and severity increase; the observer-alignment analysis (Figure 12) confirms this misalignment quantitatively.
Figure 12: Histogram of the alignment angle $\theta=\arccos|\hat{s}\cdot\hat{e}_{\text{spatial}}|$ between the optimizer boost direction $\hat{s}$ (unit spatial 3-vector of the worst-case observer’s velocity) and the spatial part $\hat{e}_{\text{spatial}}$ of the DEC-worst eigenvector of $T^{a}{}_{b}$ (the eigenvector whose principal pressure $|p_{i}|$ most exceeds $\rho$), for the Rodal metric ($50^{3}$ grid, $\zeta_{\max}=5$; see Table 1). Both vectors are normalized in the Euclidean (coordinate) norm; for the Rodal metric ($\gamma_{ij}=\delta_{ij}$) this coincides with the spatial-metric norm, so the angles are coordinate-invariant. The absolute value folds the eigenvector sign ambiguity into $0^{\circ},90^{\circ}]$. Each panel shows $\sim\!108\,000$ Type I DEC-violating points at the indicated $v_{s}$. Medians range from $76^{\circ}$ ($v_{s}=0.1$) to $85^{\circ}$ ($v_{s}=0.99$); across all four velocities the combined median is $82^{\circ}$ (near-orthogonal), confirming that the optimizer boost direction is far from the eigenvector prediction and that single-frame analysis misses violations because the Eulerian frame does not probe these directions. Only 0.20.21.21.2% of points have alignment below $10^{\circ}$.
@@
Since the vast majority of grid points classify as Hawking–Ellis Type I (Table 8), the eigenvalue criteria $\rho\geq 0$, $\rho+p_{i}\geq 0$, and $\rho\geq|p_{i}|$ are the primary determinants of energy condition satisfaction or violation at most points. For Type I, these algebraic slacks provide exact certification: a Type I point satisfies an energy condition if and only if the corresponding eigenvalue inequality holds, regardless of observer, and the result is $\zeta_{\max}$-independent. The observer optimization serves two additional roles: it identifies the worst-case observer (boost direction and rapidity), and it provides a cross-check against the algebraic result. For non-Type I stress-energy tensors, which do not admit a diagonal eigenvalue representation, the optimization becomes the sole method for evaluating observer-dependent conditions, and the algebraic shortcut does not apply. At such points, the reported WEC/SEC/DEC margins are $\zeta_{\max}$-capped lower bounds: a negative value certifies a genuine violation, but a positive value indicates only that no violation was found within the rapidity cap, not that the condition is satisfied for all observers. We validate this pathway with a null-dust (Type II) benchmark in the test suite (see Section 8).
@@
The observer optimizer with $N_{\text{starts}}=8$ random restarts is not guaranteed to find the global minimum; for highly non-convex objectives, some local minima may be missed. The soft rapidity cap $\zeta_{\max}\tanh(|\mathbf{w}|/\zeta_{\max})$ attenuates gradients near the rapidity bound, which could impede finding near-lightlike worst-case observers. Concretely, the $\tanh$ map has derivative $\mathrm{sech}^{2}(|\mathbf{w}|/\zeta_{\max})$, which decays exponentially as $|\mathbf{w}|\to\infty$; an optimizer step near the rapidity ceiling therefore receives a vanishingly small gradient signal, and worst-case observers that require $\zeta$ close to $\zeta_{\max}$ may not be found to full precision. Alternative smooth caps with slower gradient decay, such as $\zeta_{\max}\,\frac{2}{\pi}\arctan(|\mathbf{w}|)$ (polynomial $1/(1{+}w^{2})$ decay) or $\zeta_{\max}\,\frac{|\mathbf{w}|}{1+|\mathbf{w}|}$ (rational decay), would reduce this failure mode; a constrained optimizer (projected gradient or L-BFGS-B with hard bounds) would eliminate it entirely. Both alternatives are planned for a future release. The boost-vector normalization uses the smooth approximation $|\mathbf{w}|_{\varepsilon}=\sqrt{\mathbf{w}\cdot\mathbf{w}+\varepsilon^{2}}$ (with $\varepsilon=10^{-12}$) to remain $C^{\infty}$ at $\mathbf{w}=0$, avoiding discontinuous gradients that could confuse quasi-Newton Hessian updates near the Eulerian point. The ablation study in Section 7.2 confirms that margins and missed-violation fractions are effectively converged by $N_{\text{starts}}=4$, with $N_{\text{starts}}=8$ providing a safety margin.
Since most warp metrics in our test suite are predominantly Hawking–Ellis Type I, the observer optimizer is not extensively demonstrated on a regime where it is strictly necessary for violation detection (algebraic checks suffice for Type I). To validate the non-Type I pathway, we include a null-dust benchmark (Type II by construction, $T_{ab}=\Phi^{2}k_{a}k_{b}$) in the test suite, verifying that the classifier correctly identifies the null eigenvector structure and the optimizer recovers the expected NEC saturation ($T_{ab}k^{a}k^{b}=0$ to $10^{-12}$ residual) and non-negative WEC ($T_{ab}u^{a}u^{b}\geq 0$) for all tested observers; full numerical results are in the repository test suite. Full Richardson extrapolation (Section 7) was performed for the Alcubierre metric; the three-point data is non-monotone, so the assumed second-order extrapolation confirms resolution stability rather than a measured convergence rate. Rodal and WarpShell were assessed via resolution stability checks (missed-fraction and margin variation across $25^{3}$, $50^{3}$, and $100^{3}$ grids) but were not subjected to formal Richardson extrapolation, so the reported violation fractions for those metrics carry additional resolution uncertainty. Finally, the pointwise energy conditions assessed here are classical assumptions used in singularity and positive-mass theorems; they are known to be violated in semiclassical settings (e.g. Casimir effect, Hawking radiation flux). Our results are therefore best read as classical diagnostics of the metric geometry rather than universal physicality criteria.
@@
  • Shell boundaries ($R_{1}=0.5$, $R_{2}=1.0$): centers of the Hermite-smoothstep shell indicator ($C^{2}$-smooth) that controls the spatial-metric conformal factor.
@@ 9 Conclusion
    Violation severity exceeds single-frame estimates. For the Alcubierre and Natário metrics, the baseline analysis detects all NEC violation points; Lentz shows zero missed NEC/WEC points (nonzero SEC misses exist) and is severely under-resolved ($\ll\!1$ cell; this result is a lower bound). Van Den Broeck exhibits small residual NEC/WEC misses ($0.1\%$$0.4\%$ at $v_{s}=0.5$; Figures 34), and SEC misses persist (up to $1.2\%$ for Van Den Broeck, conditional miss rate $f_{\text{miss}|\text{viol}}=17.5\%$). The $\zeta_{\max}$-capped WEC severity at $\zeta_{\max}=5$ is much larger (e.g. Alcubierre WEC: $\sim\!90{,}000\times$; this ratio is $\zeta_{\max}$-dependent and grows as $\sim\!e^{2\zeta_{\max}}$). A rapidity-cap sensitivity study (Table 9) confirms that the robust WEC margin scales as $\sim\!e^{2\zeta_{\max}}$, while the NEC margin and violation-point set are cap-independent.
@@
    Tidal forces at bubble walls are significant. Geodesic deviation through the Alcubierre bubble shows sharp tidal eigenvalue spikes at the bubble wall, consistent with the expectation that steeper walls increase curvature gradients. The peak blueshift ratio at the bubble wall reaches  ${\approx}\,7.1$ at $v_{s}=0.99$, compared to  ${\approx}\,1.005$ at $v_{s}=0.1$.
@@
The author acknowledges support from the VinUni Center for Environmental Intelligence (CEI). The author thanks J. Rodal for clarifications regarding frame conventions, the $r_{s}\to 0$ limit of the irrotational angular profile, and the Cartesian form of the shift vector. The author declares no competing interests.
@@ Appendix D Metric definitions
Note on conventions (Rodal).
Ref. [20] presents the irrotational construction in a Natário-style “bubble-at-rest” convention, with radial profiles $f_{\rm paper}(0)=g_{\rm paper}(0)=0$ and $f_{\rm paper}(\infty)=g_{\rm paper}(\infty)=1$. We standardize to a lab-frame convention with vanishing shift at infinity:
$F(r_{s})=1-f_{\rm paper}(r_{s}),\qquad G(r_{s})=1-g_{\rm paper}(r_{s}),$ (32)
so that $F(0)=G(0)=1$ and $F(\infty)=G(\infty)=0$. Subtracting a uniform asymptotic translation preserves irrotationality (a uniform translation field is curl-free). We use the $\beta$” ADM convention $ds^{2}=-\alpha^{2}\,dt^{2}+\gamma_{ij}(dx^{i}+\beta^{i}\,dt)(dx^{j}+\beta^{j}\,dt)$; Ref. [20] uses $\beta$”, so component-level comparison requires the corresponding sign flip.
Removable $r_{s}\to 0$ form in $G(r_{s})$.
Although equation (31) contains an apparent $1/r_{s}$ factor, this is a removable $0/0$ form arising from the spherical-basis representation. Defining $\Delta(r_{s})\equiv\ln\cosh\bigl[\sigma(r_{s}{-}R)\bigr]-\ln\cosh\bigl[\sigma(r_{s}{+}R)\bigr]$, one finds $\Delta^{\prime}(0)=-2\sigma\tanh(\sigma R)$, whence $\lim_{r_{s}\to 0}G(r_{s})=1$. The irrotational construction is therefore regular at the bubble center; the $r_{s}\mapsto\sqrt{r_{s}^{2}+\varepsilon^{2}}$ substitution used in our implementation is a numerical stability device for automatic differentiation, not a physical repair.
Cartesian shift components.
In Cartesian coordinates, equations (30) become
$\boldsymbol{\beta}=-v_{s}\bigl[G(r_{s})\,\hat{\boldsymbol{x}}+\bigl(F(r_{s})-G(r_{s})\bigr)\,n_{x}\,\boldsymbol{n}\bigr],$ (33)
where $\boldsymbol{n}=(\Delta x,\,y,\,z)/r_{s}$ is the radial unit vector from the bubble center and $n_{x}=\Delta x/r_{s}$. Since $F(0)=G(0)=1$, the factor $(F-G)\to 0$ at $r_{s}=0$ and the shift reduces to $\boldsymbol{\beta}(0)=-v_{s}\,\hat{\boldsymbol{x}}$, confirming manifest regularity without coordinate patches.
@@
WarpShell regularization.
With $r_{\mathrm{S}}=5$ and shell radii $R_{1}=0.5$, $R_{2}=1$, the ratio $r_{\mathrm{S}}/r>1$ throughout the shell region, placing the geometry formally inside the Schwarzschild horizon where $\sqrt{1-r_{\mathrm{S}}/r}$ is imaginary in real arithmetic. The implementation applies a smooth clamp $r_{\mathrm{S}}/r\mapsto\min(r_{\mathrm{S}}/r,\;1-\epsilon)$ with $\epsilon=10^{-12}$ before evaluating the lapse and radial metric factor. This regularization produces a real, positive-definite spatial metric at the cost of replacing the Schwarzschild interior with a clamped approximation (effective lapse $\alpha_{\text{shell}}\approx\sqrt{\epsilon}\sim 10^{-6}$, radial stretch $\gamma_{rr}\sim 1/\epsilon\sim 10^{12}$ in the shell). In addition, the lapse is floored at $10^{-12}$ to ensure positivity. The resulting geometry is not a faithful Schwarzschild shell but rather a regularized toy metric designed as a numerical stress test for the classifier, optimizer, and convergence diagnostics under extreme curvature scales ($\|T^{a}{}_{b}\|\sim\mathcal{O}(10^{25})$ at $50^{3}$ resolution). Reported WarpShell energy condition results should be understood as properties of this regularized implementation, not predictions for a physically realizable warp shell spacetime. A physically consistent parameter choice (e.g. $r_{\mathrm{S}}<R_{1}$) would avoid the clamping entirely; we retain the current parameters to exercise extreme-curvature code paths.